OpenStudy (anonymous):

5/7 express as a decimal

OpenStudy (anonymous):

just the same... divide.

OpenStudy (anonymous):

5/7 *10/10?

OpenStudy (anonymous):

100/ 7 = 14.28 14.28*5 = 71.42

OpenStudy (anonymous):

Or that lol MAN! i learn everything so different :P

OpenStudy (anonymous):

1.4?

OpenStudy (anonymous):

\[\frac{5}{7} = 0.\dot{7}1428\dot{5} \]

OpenStudy (anonymous):

Haha looks like i dont learn things so different (:

OpenStudy (anonymous):

waste ur time creep

OpenStudy (anonymous):

I'm not letting someone like YOU effect me. Buh bye now!! (: